The first time I saw Linux I was unimpressed, to say the least. But as time went on the potential for hyperproductivity became obvious. The time had come to go — I waved goodbye to Bill Gates and began my transition.

The first time I saw Linux I yawned. But as time went on the potential

for hyperproductivity became ever more obvious, especially because of the LyX
software. The time had come to go — I waved goodbye to Bill Gates and began
my transition.

The first time I saw Linux I yawned. But as time went on the potential for hyperproductivity became ever more obvious, especially because of the LYX software. The time had come to go — I waved goodbye to Bill Gates and began my transition.

The first time I saw Linux I yawned. But as time went on the potential

for hyperproductivity became ever more obvious, especially because of the LyX
software. The time had come to go — I waved goodbye to Bill Gates and began

my transition.But as time went on the potential for hyperproductivity became ever more obvious, especially because of the LYX software. The time had come to go — I waved goodbye to Bill Gates and began my transition.

The first time I saw Linux I yawned. But as time went on the potential for hyperproductivity became ever more obvious, especially because of the LYX software. The time had come to go — I waved goodbye to Bill Gates and began my transition.

The first time I saw Linux I yawned. But as time went on the potential for hyperproductivity became ever more obvious, especially because of the LYX software. The time had come to go — I waved goodbye to Bill Gates and began my transition.

The first time I saw Linux I yawned. But as time went on the potential for hyperproductivity became ever more obvious, especially because of the LYX software. The time had come to go — I waved goodbye to Bill Gates and began my transition.

The first time I saw Linux I yawned. But as time went on the potential for hyperproductivity became ever more obvious, especially because of the LYX software. The time had come to go — I waved goodbye to Bill Gates and began my transition.

The first time I saw Linux I yawned. But as time went on the potential for hyperproductivity became ever more obvious, especially because of the LYX software. The time had come to go — I waved goodbye to Bill Gates and began my transition.

The first time I saw Linux I yawned. But as time went on the potential for hyperproductivity became ever more obvious, especially because of the LYX software. The time had come to go — I waved goodbye to Bill Gates and began my transition.

The first time I saw Linux I yawned. But as time went on the potential for hyperproductivity became ever more obvious, especially because of the LYX software. The time had come to go — I waved goodbye to Bill Gates and began my transition.

The first time I saw Linux I yawned. But as time went on the potential for hyperproductivity became ever more obvious, especially because of the LYX software. The time had come to go — I waved goodbye to Bill Gates and began my transition.

The first time I saw Linux I yawned. But as time went on the potential for hyperproductivity became ever more obvious, especially because of the LYX software. The time had come to go — I waved goodbye to Bill Gates and began my transition.

The first time I saw Linux I yawned. But as time went on the potential for hyperproductivity became ever more obvious, especially because of the LYX software. The time had come to go — I waved goodbye to Bill Gates and began my transition.